OpenStudy (anonymous):

Can someone remind me how to write an equation for a line parallel to an equation and containing a certain point?

OpenStudy (anonymous):

same slope

OpenStudy (anonymous):

eg y = 3x + 1 is parallel to y= 3x - 4

OpenStudy (anonymous):

What do you do to make the point show up on the line?

OpenStudy (anonymous):

like... 3x+y=5, and you need an equation with the same slope (3x) that also contains the point (-6,4)

OpenStudy (anonymous):

use the general form y-y1 = m(x-x1) where m= slope and the point is x1,y1 3x + y = 5 rearrange to slope/intercept form: y = -3x + 5 - the slope is -3 so plug in (-6,4): y-4 = -3(x - -6) y -4 = -3x -18 y = -3x - 14 is the required equn.